# Hybrid Reinforcement:When Reward Is Sparse, It’s Better to Be Dense

### v1

@@ 4.2 Main results
Table 2: The results of Qwen-3-4B-Base on both verifiable and hard-to-verify reasoning tasks. The first block reports results on verifiable tasks (MATH500, AMC, Minerva, Olympiad; with average), while the second block presents results on hard-to-verify tasks (HVM, TBR). We compare our approach—which mixes two reward signals—with two baselines: AceMath-7B-RM (a reward model) and math_verify (library), which uses a 0/1 rule as the reward.
Verifiable tasks Hard-to-verify tasks
@@
math_verify (library) 82.3 61.3 44.0 45.5 58.3 61.0 53.1 57.1
@@
AceMath-7B-RM 79.6 58.8 39.9 42.1 54.1 59.2 48.2 53.7
math_verify (library) 81.3 61.3 38.0 43.9 42.6 58.4 50.0 54.2
@@
AceMath-7B-RM 79.6 58.8 39.9 42.1 55.0 58.4 49.6 54.0
math_verify (library) 81.3 61.3 38.0 43.9 56.1 62.4 55.3 58.9
@@
Hybrid reward consistently improves performance across all regimes.
@@
Table 2 shows that HERO outperforms both reward-model-only and rule-based verifier baselines across all three regimes—verifiable, hard-to-verify, and mixed. On verifiable tasks, HERO achieves an average of 62.0, exceeding RM-only (56.4) and rule-based training (58.3). The improvement comes from stratified normalization, which allows hybrid training to exploit both positive and negative correctness groups: while verifier-only training collapses all-correct or all-incorrect batches (yielding zero relative advantage), HERO preserves meaningful gradients within each group through dense intra-group rewards. The benefit becomes most pronounced on hard-to-verify tasks, where symbolic verifiers are brittle and reward models are noisy. Here, HERO attains 56.8, surpassing RM-only (54.1) by +2.7 points and rule-based verifiers (42.6) by a large +14.2 points. This gain stems from anchoring continuous reward-model scores within verifier correctness groups, which prevents reward drift and ensures stable learning even when binary labels saturate. By combining the verifier’s precision with the reward model’s smooth discrimination, HERO can leverage partially correct reasoning paths that rule-based systems would discard, thereby improving both stability and coverage. In the mixed regime, which combines verifiable and open-ended samples, HERO again achieves the best average (58.8), outperforming RM-only (55.0) and verifier-only (56.1). Overall, hybrid reward learning delivers consistent improvements across all settings while retaining the efficiency of reward-model supervision, demonstrating that structured reward integration is critical for reasoning tasks that go beyond strict verifiability. The advantage becomes even clearer on hard-to-verify evaluations (HVM, TBR), where symbolic verifiers fail to capture partial correctness and reward models are prone to drift. Here, HERO attains 66.3 when trained on easy-to-verify data, outperforming RM-only (54.6) by +11.7 points and rule-based training (57.1) by +9.2 points. When trained on hard-to-verify samples, it still leads with 56.5 compared to 53.7 (RM-only) and 54.2 (rule-based), and under mixed training, it reaches 64.1, surpassing both baselines by large margins. These results highlight that hybrid reward design generalizes robustly across both verifiable and hard-to-verify tasks, yielding stable improvements regardless of whether evaluation relies on symbolic checking or model judgment.
Hybrid reward generalizes across backbones.
@@
Hybrid reward learning consistently scales across models with different capacities and initialization quality. On Qwen3-4B (Table 2), which already shows strong supervised performance, HERO continues to yield clear improvements across all evaluation regimes, demonstrating that hybrid optimization remains effective even for well-trained reasoning models. More strikingly, on OctoThinker-8B (Table 3), which starts from a much weaker baseline of 16.9 on verifiable and 23.6 on hard-to-verify tasks, HERO achieves substantial absolute and relative gains. When trained on easy-to-verify samples, it reaches 40.1 on verifiable and 32.6 on hard-to-verify evaluations, surpassing both the reward-model baseline (38.1/28.7) and the rule-based verifier (37.6/30.3). Training on hard-to-verify samples yields similar improvements, achieving 41.0/34.6 compared to 35.4/30.9 (RM-only) and 34.6/27.8 (verifier-only). In the mixed regime, it maintains the highest averages of 40.2 and 33.2, outperforming all baselines by 4–6 points. These results show that hybrid reward design generalizes robustly across model scales—preserving the verifier’s stability for stronger models like Qwen3-4B while bringing large relative gains to weaker ones such as OctoThinker-8B.
@@
Table 3: The results of OctoThinker-8B-Hybrid-Base on both verifiable and hard-to-verify reasoning tasks. We report pass@1 averaged over 8 seeds. The first block shows results on verifiable tasks (MATH500, AMC, Minerva, Olympiad; with Avg.), and the second block shows results on hard-to-verify tasks (HVM, TBR).
Verifiable tasks Hard-to-verify tasks
@@
math_verify (library) 60.1 39.4 26.7 24.1 37.6 31.6 28.9 30.3
@@
math_verify (library) 60.0 29.7 23.9 24.8 34.6 28.8 26.7 27.8
@@
math_verify (library) 59.3 33.7 24.7 24.0 35.4 27.6 28.7 28.2
@@
Symbolic verifiers, while offering correctness guarantees, exhibit severe limitations when applied to open-ended or format-sensitive reasoning problems. As shown in Table 2, on Qwen3-4B, verifier-only training achieves merely 42.6, which is not only lower than HERO (56.5) and RM-only (59.2) but even below the cold-start SFT baseline (47.1). A similar or even stronger degradation appears in OctoThinker-8B (Table 3), where verifier-only supervision yields only 34.6 on verifiable and 27.8 on hard-to-verify evaluations—substantially behind HERO (41.0/34.6). This performance gap widens for weaker models because their outputs are less diverse, making binary verifiers more prone to collapse when all rollouts receive identical 0 or 1 labels. Under such conditions, group-relative objectives such as GRPO produce zero gradients, halting policy improvement. By contrast, HERO introduces dense intra-group variance through reward-model refinement, maintaining effective gradient flow even in saturation regimes. This enables the policy to distinguish between partially correct and entirely incorrect solutions—an essential property for reasoning tasks where correctness lies on a continuum and purely symbolic supervision fails to provide meaningful feedback.
@@ 4.3 Additional ablations
We found that dense negative rewards play a more critical role in stabilizing training and improving learning efficiency than positive samples. While positives signal correctness, negatives provide richer supervision by penalizing diverse reasoning errors. Notably, using only negative rewards boosts performance on verifiable tasks from 59.4 to 61.4, and even more on hard-to-verify tasks from 62.2 to 68.4. This demonstrates that well-calibrated negative ranges are essential: they provide broader feedback, enabling the model to detect subtle errors and generalize better on complex cases.
@@
We conducted ablation studies to investigate the impact of varying reward ranges on model performance, as shown in Figure 2(b). For verifiable tasks, smaller reward ranges (e.g., $\alpha=0.05$) yielded the best results, as the rule-based verifier’s precision benefits from a tighter range that minimizes noise and maintains stability. Expanding the range beyond this threshold led to diminishing returns and increased noise. In contrast, for mixed tasks, where many samples fail the rule-based verifier, the learned reward model plays a larger role. Here, larger reward ranges (e.g., $\alpha=0.1$ or $\alpha=0.2$) provided richer signals, allowing the model to learn more effectively from harder tasks. However, expanding the range beyond a certain point caused a slight performance drop due to overfitting or excessive noise. Overall, careful tuning of the reward range, particularly for the negative rewards, is crucial to balancing stability and performance, depending on the task type.
Variance-aware reweighting in HERO improves model’s reasoning ability.
@@
Methods Verifiable Hard-to-verify
@@
We evaluated variance-aware reweighting based on reward-model score variance, which emphasizes ambiguous, high-variance prompts while down-weighting trivial ones to reduce overfitting. This dynamic adjustment yields consistent gains, particularly on hard-to-verify tasks where dense signals are most informative. As shown in Table 4, reweighting improves accuracy on both verifiable and hard-to-verify benchmarks, with larger gains in the latter (+3.8), confirming that focusing capacity on uncertain samples leads to more robust and generalizable improvements.
@@
Table 5: Impact of reward model size: a larger RM provides no remarkable gain over the smaller RM.
Reward model Verifiable Hard-to-verify
@@
A natural question is whether stronger supervision requires scaling up the reward model itself. To test this, we compare HERO with AceMath-RM-7B against a much larger variant, AceMath-RM-72B. As shown in Table 5, the larger reward model does not provide meaningful gains: while it slightly improves performance on verifiable tasks (62.8 vs. 62.0), it underperforms on hard-to-verify tasks (71.4 vs. 73.2). This result suggests that the benefits of our framework stem primarily from hybrid design—stratified normalization and variance-aware weighting—rather than from simply scaling the reward model. In practice, this means HERO can achieve strong results with compact reward models, offering better efficiency and deployability without sacrificing accuracy.
@@ 5 Related Work
Reinforcement learning from verifiable rewards (RLVR) leverages deterministic correctness checks—such as passing unit tests or matching reference answers—to enhance model learning (Shao et al., 2024). Early program synthesis work demonstrated that agent-generated trajectories validated against ground truth outperform supervised approaches (Bunel et al., 2018; Chen et al., 2021). In the context of LLMs, rule-based verification plays a crucial role in filtering, providing training signals, and supporting benchmark evaluations (Xiong et al., 2025; Yu et al., 2025; Shao et al., 2024). Recent extensions include: outcome-driven RL (GRPO) for grounding and citation fidelity in QA tasks (Sim et al., 2025); rubric-anchored RL, which introduces structured rubrics for open-ended response evaluation (Huang et al., 2025b); verifier-free RL strategies like VeriFree, which bypass explicit checking while achieving performance on par with verifier-based methods (Zhou et al., 2025); and cross-domain RLVR, which employs LLM-derived scoring for domains lacking clear reference answers (Su et al., 2025). Despite these advancements, rule-based methods still struggle with semantically correct but textually divergent outputs, motivating the use of model-based verifiers (Chen et al., 2025a; Ma et al., 2025; Huang et al., 2025a; Xu et al., 2025). However, the coverage of LLM-based verifiers remains limited for the generalization (Li et al., 2025), and the rewards they provide are still sparse, often consisting of binary labels. In contrast to previous work, we propose a hybrid approach that combines rule-based verification with continuous, dense reward signals from learned models, allowing us to maintain the stability of verifiers while addressing their sparsity. By anchoring dense signals to symbolic correctness and introducing a variance-aware weighting mechanism, our method enables more informative, stable, and sample-efficient learning on both verifiable and hard-to-verify tasks.
@@
As the reasoning capabilities of large language models (LLMs) have reached new heights, increasingly challenging reasoning benchmarks have been proposed (Phan et al., 2025; Zhang et al., 2025a). These problems often involve complex outputs, such as natural language representations and intricate mathematical or physical formulas. In such cases, rule-based verification methods, while effective for well-defined problems, struggle to capture the nuances of these tasks. Recent work has focused on the use of LLMs as judges, where LLMs assess the quality of generated responses (Chen et al., 2025a; Ma et al., 2025; Huang et al., 2025a; Xu et al., 2025; Li et al., 2025), enabling more nuanced evaluations. However, despite its conceptual simplicity, LLM-as-judge may not always produce reliable assessments for domain-specific or long-form data. Some recent work proposes going beyond binary labels from verifiers for hard-to-verify tasks. For example, Gurung & Lapata (2025) applies reasoning traces in Next-Chapter Prediction (NCP) for long-form story generation via likelihood estimation, while Tang et al. (2025) uses Jensen’s evidence lower bound to treat chain-of-thought reasoning steps as latent variables in the generative process. They directly get rid of the verifier component. In contrast, our work retains the use of verifiable rewards but enhances supervision through the introduction of a reward model.
@@ 6 Conclusion
We introduced HERO, which anchors reward-model signals to verifier-defined correctness via stratified normalization and emphasizes informative prompts with variance-aware weighting. This hybrid design preserves the stability of verifiers while supplying dense, trajectory-sensitive feedback, mitigating gradient sparsity and RM-only drift. Empirically, HERO consistently outperforms RM-only and verifier-only baselines across verifiable, hard-to-verify, and mixed regimes and across backbones. Future work includes stronger difficulty estimators, process-level rewards, and extension beyond math reasoning.

### v3

@@ 4.2 Main results
Table 2: Performance of HERO trained with Qwen3-4B-Base on both easy-to-verify and hard-to-verify reasoning tasks. The first block reports results on verifiable tasks (MATH500, AMC, Minerva, Olympiad; with average), while the second block presents results on hard-to-verify tasks (HVM, TBR). We compare our approach HERO —which combines two reward signals—with two baselines corresponding to the two signals: AceMath-7B-RM (a reward model) and math_verify (verl), which uses a 0/1 rule as the reward.
Easy-to-verify tasks Hard-to-verify tasks
@@
math_verify (verl) 82.3 61.3 44.0 45.5 58.3 61.0 53.1 57.1
@@
AceMath-7B-RM 79.6 58.8 39.9 42.1 55.1 59.2 48.2 53.7
math_verify (verl) 76.2 46.6 28.7 38.2 47.4 58.4 50.0 54.2
@@
AceMath-7B-RM 79.6 58.8 39.9 42.1 55.1 58.4 49.6 54.0
math_verify (verl) 81.3 61.3 38.0 43.9 56.1 62.4 55.3 58.9
@@
Performance of HERO on the Qwen-based Model
@@
Table 2 shows that HERO consistently outperforms both the reward-model-only and rule-based verifier baselines across all three training data settings: (1) easy-to-verify data, (2) hard-to-verify data, and (3) a mixture of the two. For each training setting, we evaluate on four datasets where the targets are easy-to-verify tasks, as well as two datasets where the targets are hard-to-verify tasks. When trained on easy-to-verify data and evaluated on four easy-to-verify test sets, HERO achieves an average score of 62.0, outperforming both RM-only (56.4) and rule-based training (58.3). We attribute this improvement to our stratified normalization, which allows hybrid training to exploit both positive and negative correctness groups: while verifier-only training collapses all-correct or all-incorrect batches (yielding zero relative advantage), HERO preserves meaningful gradients within each group through dense intra-group rewards. The advantage of our approach becomes even more pronounced when training on hard-to-verify samples, where rule-based verifiers are brittle and reward models tend to be noisy. Here, HERO attains 56.8, surpassing RM-only (55.1) by +1.7 points and rule-based verifiers (47.4) by a substantial +9.4 points. This improvement is due to anchoring continuous reward-model scores within verifier correctness groups, which prevents reward drift and ensures stable learning even when binary labels saturate. By combining the precision of rule-based verifiers with the smooth discrimination of reward models, HERO is able to leverage partially correct reasoning paths that would otherwise be discarded by rule-based systems, thereby improving both stability and coverage. When trained on mixed data, which combines easy-to-verify and hard-to-verify samples, HERO again achieves the average (58.8), outperforming RM-only (55.1) and rule-based verifier (56.1) on verifiable tasks.
The advantage becomes even clearer on hard-to-verify evaluations (HVM, TBR), where rule-based verifiers fail to capture partial correctness and reward models are prone to drift. Here, HERO attains 66.3 when trained on easy-to-verify data, outperforming RM-only (54.6) by +11.7 points and rule-based training (57.1) by +9.2 points. When trained on hard-to-verify samples, it still leads with 56.5 compared to 53.7 (RM-only) and 54.2 (rule-based). Under mixed training, HERO reaches 64.1, surpassing both baselines by large margins on hard-to-verify tasks. These results highlight that hybrid reward design generalizes robustly across both verifiable and hard-to-verify tasks, yielding stable improvements regardless of whether evaluation relies on symbolic checking or model judgment. Overall, hybrid reward learning delivers consistent improvements across all settings, demonstrating that structured reward integration is critical for reasoning tasks that go beyond strict verifiability.
Performance of HERO on the OctoThinker-based Model
@@
On Qwen3-4B-Base (Table 2), which already shows strong performance, HERO consistently delivers clear improvements across all evaluation settings. On OctoThinker-8B-Hybrid-Base (Table 3), which starts from a much weaker baseline of 16.9 on verifiable and 23.6 on hard-to-verify tasks, HERO achieves substantial absolute and relative gains. When trained on easy-to-verify samples, it reaches 40.1 on verifiable and 32.6 on hard-to-verify evaluations, surpassing both the reward-model baseline (38.1/28.7) and the rule-based verifier (37.6/30.3). Training on hard-to-verify samples yields similar improvements, achieving 41.0/34.6 compared to 35.4/30.9 (RM-only) and 34.6/27.8 (verifier-only). Training on mixed training samples, it maintains the highest averages of 40.2 and 33.2, outperforming all baselines by 4–6 points. These results show that hybrid reward design generalizes robustly across model scales—preserving the verifier’s stability for stronger models like Qwen3-4B-Base while bringing large relative gains to weaker ones such as OctoThinker-8B-Hybrid-Base.
@@
Table 3: Performance of HERO trained with OctoThinker-8B-Hybrid-Base on both easy-to-verify and hard-to-verify reasoning tasks. The first block shows results on four easy-to-verify tasks, which reported pass@1 averaged over 8 seeds. The second block show results on two hard-to-verify testsets, which reported GPT4.1 judges results.
Easy-to-verify tasks Hard-to-verify tasks
@@
math_verify (verl) 60.1 39.4 26.7 24.1 37.6 31.6 28.9 30.3
@@
math_verify (verl) 60.0 29.7 23.9 24.8 34.6 28.8 26.7 27.8
@@
math_verify (verl) 59.3 33.7 24.7 24.0 35.4 27.6 28.7 28.2
@@
Symbolic verifiers, while offering correctness guarantees, exhibit severe limitations when applied to open-ended or format-sensitive reasoning problems. As shown in Table 2, on Qwen3-4B-Base, verifier-only training achieves merely 47.4 for average on hard-to-verify tasks, which is not only lower than HERO (56.5) and RM-only (53.7) but even below the cold-start SFT baseline (47.1). A similar or even stronger degradation appears in OctoThinker-8B-Hybrid-Base (Table 3), where verifier-only supervision yields only 34.6 on verifiable and 27.8 on hard-to-verify evaluations—substantially behind HERO (41.0/34.6). This performance gap widens for weaker models because their outputs are less diverse, making binary verifiers more prone to collapse when all rollouts receive identical 0 or 1 labels. Under such conditions, group-relative objectives such as GRPO produce zero gradients, halting policy improvement. By contrast, HERO introduces dense intra-group variance through reward-model refinement, maintaining effective gradient flow even in saturation regimes. This enables the policy to distinguish between partially correct and entirely incorrect solutions—an essential property for reasoning tasks where correctness lies on a continuum and purely symbolic supervision fails to provide meaningful feedback.
@@ 4.3 Additional ablations
We evaluate the role of dense negative and positive reward on the setting of training with easy-to-verify samples based on Qwen-4B-Base. We found dense reward in the negative group play a more criticak role in stabilizing training and improving learning efficiency than dense reward in the positive group devided by HERO.. While positives signal correctness, negatives provide richer supervision by penalizing diverse reasoning errors. Notably, dense negative rewards but maintaining sparse verifier positive rewards boosts performance on verifiable tasks from 59.4 to 61.4, and even more on hard-to-verify tasks from 62.2 to 68.4 (see Figure 2). This demonstrates that well-calibrated negative ranges are essential: they provide broader feedback, enabling the model to detect subtle errors and generalize better on complex cases.
@@
We conducted ablation studies to investigate the impact of varying reward ranges on model performance by training with easy-to-verify samples based on Qwen-4B-Base, as shown in Figure 2(b). For verifiable tasks, smaller reward ranges (e.g., $\alpha=0.05$) yielded the best results, as the rule-based verifier’s precision benefits from a tighter range that minimizes noise and maintains stability. Expanding the range beyond this threshold led to diminishing returns and increased noise. In contrast, for mixed tasks, where many samples fail the rule-based verifier, the learned reward model plays a larger role. Here, larger reward ranges (e.g., $\alpha=0.1$ or $\alpha=0.2$) provided richer signals, allowing the model to learn more effectively from harder tasks. However, expanding the range beyond a certain point caused a slight performance drop due to overfitting or excessive noise. Overall, careful tuning of the reward range, particularly for the negative rewards, is crucial to balancing stability and performance, depending on the task type.
Variance-aware reweighting in HERO improves the model’s reasoning ability.
@@
Methods Easy-to-verify Hard-to-verify
@@
We evaluated variance-aware reweighting based on reward-model score variance, which emphasizes ambiguous, high-variance prompts while down-weighting trivial ones to reduce overfitting with the setting of training with easy-to-verify samples based on Qwen-4B-Base. This dynamic adjustment yields consistent gains, particularly on hard-to-verify tasks where dense signals are most informative. As shown in Table 4, reweighting improves accuracy on both verifiable and hard-to-verify benchmarks, with larger gains in the latter (+3.8), confirming that focusing capacity on uncertain samples leads to more robust and generalizable improvements.
@@
Table 5: Impact of reward model size: a larger RM provides in HERO no remarkable gain over the HERO with smaller RM .
Reward model Easy-to-verify Hard-to-verify
@@
A natural question is whether stronger supervision requires scaling up the reward model itself. To isolate this factor, we replace the 7B reward model in HERO with a much larger 72B reward model, keeping the verifier and all training configurations fixed. As shown in Table 5, the larger reward model yields only a marginal improvement on verifiable tasks (62.8 vs. 62.0) and even slightly underperforms on hard-to-verify tasks (71.4 vs. 73.2). This confirms that the gains of HERO primarily come from its hybrid reward formulation—through stratified normalization and variance-aware weighting—rather than from reward model scaling. Practically, this means that HERO can achieve strong results with compact reward models, offering better efficiency and deployability without sacrificing accuracy.
@@ 5 Related Work
Reinforcement learning from verifiable rewards (RLVR) leverages deterministic correctness checks—such as passing unit tests or matching reference answers—to enhance learning (Shao et al., 2024). Early program synthesis work demonstrated that agent-generated trajectories validated against ground truth outperform supervised approaches (Bunel et al., 2018; Chen et al., 2021). For LLMs, rule-based verification plays a crucial role in filtering, providing training signals, and supporting benchmark evaluations (Xiong et al., 2025; Yu et al., 2025; Shao et al., 2024). Recent extensions include: outcome-driven RL (GRPO) for grounding and rubric-anchored RL which introduces structured rubrics for open-ended response evaluation (Huang et al., 2025b); verifier-free RL strategies like VeriFree, which bypass explicit checking by directly maximizing the probability of generating the reference answer while achieving performance on par with verifier-based methods (Zhou et al., 2025); and cross-domain RLVR, which employs LLM-derived scoring for domains lacking reference answers (Su et al., 2025). Despite these advancements, rule-based methods still struggle with semantically correct but textually divergent outputs, motivating the use of model-based verifiers (Chen et al., 2025a; Ma et al., 2025; Huang et al., 2025a; Xu et al., 2025). However, the coverage of LLM-based verifiers remains limited the generalization (Li et al., 2025), and the rewards they provide are sparse, often consisting of binary labels. In contrast to previous work, we propose a hybrid approach that combines rule-based verification with continuous, dense reward signals from learned models, allowing us to maintain the stability of verifiers while addressing their sparsity. By anchoring dense signals to symbolic correctness and introducing a variance-aware weighting mechanism, our method enables more informative, stable, and sample-efficient learning on both verifiable and hard-to-verify tasks.
@@
As the reasoning capabilities of LLMs have reached new heights, increasingly challenging reasoning benchmarks have been proposed (Phan et al., 2025; Zhang et al., 2025a). These problems often involve complex outputs, such as natural language representations and intricate mathematical or physical formulas. In such cases, rule-based verification methods, while effective for well-defined problems, struggle to capture the nuances of these tasks. Recent work focused on the LLMs as judges, where LLMs assess the quality of generated responses (Chen et al., 2025a; Ma et al., 2025; Huang et al., 2025a; Xu et al., 2025; Li et al., 2025), enabling more nuanced evaluations. However, despite its conceptual simplicity, LLM-as-judge may not always produce reliable assessments for domain-specific or long-form data. Some recent work proposes going beyond binary labels from verifiers for hard-to-verify tasks. For example, Gurung & Lapata (2025) applies reasoning traces in Next-Chapter Prediction (NCP) for long-form story generation via likelihood estimation, while Tang et al. (2025) uses Jensen’s evidence lower bound to treat chain-of-thought reasoning steps as latent variables in the generative process. They directly discard the verifier component. In contrast, our work retains the use of verifiable rewards but enhances supervision through the introduction of a reward model. Peng et al. (2025) directly add the verifiable correctness signals and the human preferences for agentic tasks.
@@ 6 Conclusion
We introduced HERO (Hybrid Ensemble Reward Optimization), which anchors reward-model signals to verifier-defined correctness via stratified normalization and emphasizes informative prompts with variance-aware weighting. This hybrid design preserves the stability of verifiers while supplying dense, trajectory-sensitive feedback, mitigating gradient sparsity and RM-only drift. Empirically, HERO consistently outperforms RM-only and verifier-only baselines across verifiable, hard-to-verify, and mixed regimes and across backbones. Future work includes examining stronger difficulty estimators, process-level rewards, and extensions beyond math reasoning.